OpenStudy (anonymous):

What is the geometric mean of 3 and 15

OpenStudy (anonymous):

\[\sqrt{3\times 15}=\sqrt{45}=\sqrt{9\times 5}=\sqrt{9}\sqrt{5}=3\sqrt{5}\]

OpenStudy (anonymous):

what about this one : Six is the geometric mean of 18 and what number? and thank u soo much

OpenStudy (anonymous):

2

OpenStudy (anonymous):

@amir.sat i had asked onther question but not on this one can u help me with it plzz i can post it on here

OpenStudy (anonymous):

a,b,c are geometric series if ac=b^2

OpenStudy (anonymous):

thank u both of u soo much